OpenStudy (anonymous):

f(x) -3(x-1)^2 / (x-2)(x+2)^2 What is the horizontal asymptote?

OpenStudy (anonymous):

Anyone?

OpenStudy (freckles):

what is the deg of the top and the deg of the bottom?

OpenStudy (freckles):

if the degs are the same then you take coefficients of the term with highest exponents on both top and bottom if the deg of the bottom is bigger than the deg of the top, then your horizontal asymptote is just y=0 if the deg of the top is bigger than the deg of the bottom, then you don't have a horizontal asymptote.

OpenStudy (anonymous):

So basically it's y=0?

OpenStudy (freckles):

yeah if this is the function I'm reading: \[f(x)=\frac{-3(x-1)^2}{(x-2)(x+2)^2}\] deg on top=2 deg on bot=(1+2)=3 deg on bot>deg on top